### v1

# Diagnostic Reasoning in Natural Language: Computational Model and Application

Abstract
Diagnostic reasoning is a key component of expert work in many domains. It is a hard, time-consuming activity that requires expertise, and AI research has investigated the ways automated systems can support this process. Yet, due to the complexity of natural language, the applications of AI for diagnostic reasoning to language-related tasks are lacking. To close this gap, we investigate diagnostic abductive reasoning (DAR) in the context of language-grounded tasks (NL-DAR). We propose a novel modeling framework for NL-DAR based on Pearl’s structural causal models and instantiate it in a comprehensive study of scientific paper assessment in the biomedical domain. We use the resulting dataset to investigate the human decision-making process in NL-DAR and determine the potential of LLMs to support structured decision-making over text. Our framework, open resources and tools lay the groundwork for the empirical study of collaborative diagnostic reasoning in the age of LLMs, in the scholarly domain and beyond. [^1] Code and data will be publicly released upon acceptance.
1 Introduction
Diagnostic abductive reasoning (DAR) is an important part of human decision-making in expert domains. Given observations about a system, DAR seeks to find the most plausible explanation for the observed behavior in terms of the system’s hidden states Reiter (1987). Applications of DAR are manifold: doctors use diagnostic reasoning to determine the disease based on symptoms and measurements, and programmers diagnose code to determine the faulty software component that leads to the observed runtime behavior Magnani (2023). Diagnostic reasoning is a hard expert task, and a long line of research has explored the use of AI to expedite this process (Peng and Reggia, 1990; Richens et al., 2020, i.a.). Recent applications of AI to support DAR that decompose problems of the physical world into step-wise diagnostic graphs, such as in the medical domain, show great promise Richens and Buchard (2022).
Figure 1: We use NL-DAR framework to decompose paper assessment into a set of causally connected steps, incl. reading the paper, extracting the evidence, reasoning over the evidence, and arriving at a verdict.
Assessing documents written in natural language is a key part of day-to-day knowledge work, and it involves a lot of diagnostic thinking. From essay grading to fact-checking to scholarly peer review, experts reason about a document based on specific observations (e.g., "Does the news report cite its sources?", "Does the paper report participant details?") and arrive at a conclusion ("the article might be unreliable", "the results might be biased"). Yet, for most textual assessment tasks, only the outcome (e.g., a factuality label or the review report) and not its underlying reasoning process are recorded. This leads to data scarcity that hinders the training and evaluation of natural language processing models using the existing DAR methodology. The potential of natural language DAR (NL-DAR) thus remains under-explored.
Recent breakthroughs in large language models (LLMs) create new opportunities for applying AI to diagnostic reasoning grounded in natural language. In particular, the ability to perform tasks in a zero-shot fashion and to reason over the inputs Brown et al. (2020); Kojima et al. (2022) can alleviate the current methodological hurdles that prevent the application of DAR to natural language problems. Explicit reasoning frameworks, in turn, have the potential to enhance the interpretability and transparency of LLM applications to complex diagnostic tasks, and to steer human-AI collaboration. Thus, investigating LLMs in the context of diagnostic reasoning bears great promise for both lines of research.
Yet, open questions remain. Diagnostic reasoning commonly involves multiple interconnected reasoning steps. (RQ1) How to represent the diagnostic reasoning processes for NL-DAR problems? To approach this, we introduce workflows – an application of structural causal models Pearl (1995) that decompose diagnostic reasoning into a step-wise reasoning graph starting with the text observations and leading towards an outcome. (RQ2) How to devise a workflow for a particular NL-DAR application? We introduce and test a three-step process where we construct workflows based on expert interactions and use the resulting causal models to empirically study how humans perform the task. (RQ3) How can state-of-the-art large language models support NL-DAR? To answer, we perform extensive experiments with four state-of-the-art LLMs and report on the ability of LLMs to perform individual diagnostic reasoning steps as part of an NL-DAR workflow.
Diagnostic reasoning with natural language covers a wide spectrum of problems. To illustrate the application of our framework, we report on a large-scale study in biomedical paper assessment – a task closely related to the emerging line of research in AI-assisted peer review. We enroll over $40$ researchers and students in the biomedical field and develop a practical annotation interface based on the CARE platform Zyska et al. (2023) that allows users to manually perform workflow-based assessment of texts. The resulting large-scale dataset provides an empirical basis for the study of human diagnostic reasoning and LLM assistance in the domain of scholarly peer review, and the toolkit enables applying our proposed framework to other NL-DAR problems. In total, this work contributes:
  • a formal model for natural language abductive diagnostic reasoning tasks rooted in causality theory (RQ1) and a framework for applying this model to specific NL-DAR tasks (RQ2);
  • the first expert-generated multi-modal diagnostic reasoning dataset in biomedical paper assessment, comprising of more than $4$k text-based reasoning steps on top of $22$ biomedical papers;
  • insights into human influence factors during paper assessment (RQ2) and the potential of LLM assistants in diagnostic abductive reasoning for paper assessment (RQ3).
In our experiments, we find that (RQ2) differences in prior knowledge are a central source of disagreement among humans and that a paper’s writing style has a major impact on the overall assessment. We further find that (RQ3) LLMs suffer from error propagation in the reasoning graph but that human oversight effectively mitigates this issue, showing the overall promise of the proposed approach. Our study bridges the gap between the research in causality, human-AI collaboration, and LLMs, and lays the groundwork for applying LLMs to expert tasks that involve complex diagnostic reasoning grounded in natural language.
2 Related Work
We contribute to abductive reasoning research in NLP Bhagavatula et al. (2020). Prior studies primarily focus on abduction based on formal logic converted into natural language (e.g. Young et al., 2022). In contrast, we take an application-driven approach, which involves complex, subjective expert reasoning on a real-world task. Similar to our study, Dalal et al. (2024) evaluate LLMs’ ability to infer the best explanation for an observation and provide its reasoning. Unlike this work, we impose a reasoning structure on the model to ensure high-quality explanations. Our work is related to research on explainability in NLP (e.g. Wiegreffe and Marasovic, 2021; Camburu et al., 2018). In particular, Dalvi et al. (2021) test models on constructing reasoning paths over short facts (entailment trees) as explanations for science questions. Our framework generalizes entailment trees to directed graphs that serve as explanations. Our work connects NLP with research on AI for DAR. Existing works in this area focus on medical Richens and Buchard (2022) and engineering problems Koitz-Hristov and Wotawa (2018) that rely primarily on bipartite diagnosis graphs. Our framework for DAR in natural language considers graphs of more than two layers; it adds natural language variables to the diagnosis graph and is the first to use LLMs to model the relations between diagnostic variables. Finally, we contribute to NLP for peer review support Kuznetsov et al. (2024). We complement prior work on argumentation and intentions in review reports Hua et al. (2019); Cheng et al. (2020); Kuznetsov et al. (2022); Dycke et al. (2023) by the first empirical analysis of the reasoning that underlies review writing. D’Arcy et al. (2024) and Chamoun et al. (2024) propose multi-agent LLM architectures to generate feedback on scientific papers. In contrast, we test LLMs on the reasoning steps leading towards the assessment of a paper.
3 Framework
3.1 Background: Causal Models
Causality provides a formal way to reason about cause-and-effect relationships by specifying how active changes to one variable (interventions) change the likelihood of another. One widely used type of causal models, adopted in this work, are Structural Causal Models (SCMs, Pearl (1995)). Following the definition by Pearl (2000):
Definition 1
An SCM $M$ is $(U,V,F,P_{M})$, where
  1. i.
    $U$ is a set of exogenous variables determined by factors outside of $M$. $P_{M}$ is a probability function over the domain of $U$.
  2. ii.
    $V=\{V_{1},...,V_{n}\}$ is a set of endogenous variables determined by variables in $U\cup V$.
  3. iii.
    $F=\{f_{1},...,f_{n}\}$ is a set of functions on the domains of $U_{i}\cup\textit{pa}_{i}$ to $V_{i}$, where $\textit{pa}_{i}\subseteq V\setminus V_{i}$, and $U_{i}\subseteq U$ and the entire set $F$ forms a mapping from $U$ to $V$.
This means each endogenous (i.e., modeled) variable is given by a structural equation $v_{i}=f_{i}(\textit{pa}_{i},u_{i})$ where $\textit{pa}_{i}$ is the set of parent variables from $V$ and arbitrary exogenous variables $u_{i}\subseteq U$. Exogenous variables are also called background variables that add randomness to the otherwise deterministic structural equations.
Causal Graphs
Causal graphs are an intuitive representation of SCMs where the nodes are given by $U\cup V$, and a directed edge indicates that the antecedent is input to the structural equation of the consequent (child). We read an edge from $X$ to $Y$ as X (may) cause Y.
Interventions and Counterfactuals
We notate interventions using the $do$-operator. For an intervention $do(X=x)$, we alter the SCM $M$ by replacing all occurrences of variable $X$ by $x$ in $F$. We denote the altered model as $M_{X=x}$. Counterfactuals pose what-if questions on the SCM. For a query "What value had $Y$ if $X=x$ under the circumstances $u$?" we imagine the causal effect of $X$ on $Y$ for a concrete instance counter to the observed facts. The counterfactual distribution is
$P_{M_{X=x}}(Y=y)=\sum_{\{u|u\in\widehat{U}\land Y_{x}(u)=y\}}P(u)$
where $Y_{x}(u)$ denotes the potential response of $Y$ under $M_{X=x}$. Finally, the average causal effect (ACE) of an intervention on a binary variable $X$ is the difference of expected values of $Y$ with $\textit{ACE}(X,Y)=\mathbb{E}[Y_{x=1}]-\mathbb{E}[Y_{x=0}]$. We use ACE and counterfactuals during analysis in Section 4.3.
3.2 Natural Language Diagnostic Reasoning
At the heart of diagnostic abductive reasoning lies a system model that describes the relations between components and the observable behavior Peng and Reggia (1990). We now formalize an NL-DAR problem using SCMs to specify the system model. [^2] Various formalisms have been proposed for DAR, including fuzzy logic, and probabilistic graphical models Peng and Reggia (1990); Richens et al. (2020). SCMs are strictly more powerful than these formalisms Pearl (2000).
Definition 2
A natural language diagnostic abductive reasoning problem is given by an SCM $M=(U,V,F,P_{M})$, the workflow, with
  1. i.
    inputs $I\subset V$ and $I\neq\emptyset$ in natural language.
  2. ii.
    final verdicts on the input with $T\subseteq V$ and $T\neq\emptyset$ analogous to the system behavior.
  3. iii.
    components $C\subset V$ with $C\cap I=\emptyset$, $C\cap T=\emptyset$ and $C\neq\emptyset$ forming the explanation, i.e. the steps of the workflow.
All inputs $i\in I$ are roots of the causal graph, i.e., $\textit{pa}_{i}=\emptyset$, and form input to the structural equation of at least one component, i.e., $\forall i\in I:\exists c\in C:i\in\textit{pa}_{c}$. The final verdicts have no children $\forall t\in T:\forall v\in V:t\notin\textit{pa}_{v}$. Background variables determine the level of subjectivity and noise in the task. Solving an NL-DAR problem means inferring the most likely values for all $c\in C$ given $I$ and $T$.
Less formally, a workflow describes a reasoning graph from an input text to the verdict, and the diagnosis is the answers to all intermediate steps explaining the verdict and depending on the textual input. Notably, Definition 2 deviates from DAR problems in the physical world, as the observation (verdict) does not exist on its own but arises as a result of the interpretation of the text by a reader.
3.3 Approaching an NL-DAR Problem
Our definition of NL-DAR implies a three-step approach to problems of this class.
Step 1. Designing the SCM
First, the endogenous variables that might impact the final verdicts must be defined. Following that, the SCM’s structure is inferred from observational data through causal discovery Verma and Pearl (1990) or determined through interaction with domain experts. The steps of the workflow are modeled as boolean decisions, numerical scores, or natural language answers [^3] Causality-aware text representation are a hard research problem without a general solution Veitch et al. (2020); Feder et al. (2022). In our case study (Section 4), we combine boolean and natural language answers.. This process yields the graph structure and variables of the workflow.
Step 2. Populating the SCM
Next, structural equations are estimated empirically. To do this, human annotators are guided through the workflow, starting with the input document and moving toward the output verdicts. The intermediate steps serve as the individual explanations of the human’s final verdicts. Varying human experiences and biases act as background factors $U$. If a diagnostic task is subjective, background factors induce high variability in the answers. To measure the level of noise, each step and input is annotated redundantly. Based on this data, one can estimate the structural equations using machine learning models trained on the input-output data for each workflow step.
Step 3. Analysis and automation
Once the underlying SCM is defined and populated, one can analyze it to empirically describe the human diagnostic reasoning process through causal queries on the model. The SCM can be further used as the backbone to solve new instances of specific NL-DAR problems automatically or via human-AI collaboration. To solve new instances of the NL-DAR problem, causal inference is applied, providing new inputs to the trained machine learning models of the structural equations. As we show in Section 5, LLMs naturally combine with the causal graph to solve NL-DAR problems with natural language reasoning steps, devising the SCM as a probabilistic LLM program Dohan et al. (2022).
4 Paper Assessment as NL-DAR
Figure 2: Paper assessment workflow (excerpt). Information from the paper is acquired and aggregated in intermediate reasoning steps. The overall assessment lies at the end of the graph. Each node is associated with a prompt, description, example, and output format. The arrows are causal dependencies reading as therefore.
The growing publication rates across sciences put a strain on academic quality control mechanisms like peer review, which could greatly benefit from AI assistance Kuznetsov et al. (2024). We now report on the application of the proposed framework to biomedical paper assessment cast as an NL-DAR problem. Following Definition 2, our workflow describes the process where humans extract bits of information from the paper text and come to an overall judgment (the verdict) based on a mental model of what constitutes a high-quality paper.
4.1 Designing the SCM
We derive the SCM for biomedical text assessment based on expert interviews followed by an independent validation step. Following prior work on modeling complex systems with expert-elicited SCMs Rodrigues et al. (2022); Ashdown and Petrou (2023), we derive a best-effort SCM that is agreeable with the mental paper quality model of domain experts while accounting for subjectivity via the SCM’s background variables. We separately interview two experienced biomedical researchers (postdoc and professor) to understand the criteria for assessing a scientific article in this field. We aggregate and structure their responses to form a step-wise reasoning graph (detailed below), which we revise together with the two experts. Finally, we validate the graph with three PhD students from the same field in a formative study where we ask them to assess a biomedical article following the reasoning graph. Appendix A.1.1 provides details.
The resulting workflow comprises $45$ interdependent paper quality assessment steps terminating in a final verdict on the paper quality. Figure 2 illustrates an excerpt of the workflow. Based on the expert interviews, we identify three key activities during the assessment:
  • for read steps, humans read a portion of the paper in detail.
  • for extract steps, humans distill certain information from the previously read text, e.g., listing the results implied by a figure.
  • for infer steps, humans reason over prior answers to combine or reduce information into assessments, e.g., "Are the results plausible?".
We associate each step in the graph with one of these activities, a task prompt, a name, detailed instructions, and an example (see Appendix A.1.2).
The ranges of the endogenous variables are subject to design choice. The assessment of a paper is coupled with its text and figures, requiring a multimodal representation of the inputs. Likewise, in the formative study, we identified that while some steps have boolean values (e.g., "Are the experimental results plausible given the results of related work?"), humans tend to provide detailed, unstructured answers to them. Therefore, we opt for a textual representation of endogenous variables while allowing additional boolean answers for infer nodes and permitting images on the read steps.
4.2 Populating the SCM
To estimate the structural equations of the SCM from data, we collect a dataset of human reasoning steps following the causal graph structure. We guide the participants through the workflow to assess a range of papers while redundantly annotating the same paper for the estimation of noise.
Source data.
We use a sample of openly licensed biomedical manuscripts from bioRxiv [^4] https://www.biorxiv.org/ filtered by recency (06/2021 to 06/2023) and expertise keywords and stratified by topic, world region, and publication status (number of revisions and acceptance status at a peer-reviewed venue). The resulting stratified sample of $200$ papers is further manually filtered by an expert annotator to ensure sufficient expertise and is subject to subselection based on annotator preference, resulting in $22$ papers. See Appendix A.2.1 for more details.
Tools and guidelines
The annotation study was performed in two stages, which we further denote junior and senior based on the level of the participants’ academic expertise. We use similar guidelines, setup, and annotation interfaces for both stages. We turn the graphical workflow into a sequential questionnaire, selecting a permutation of steps that follows the paper structure and adheres to the topological ordering of the graph. This ensures that at any step, the prior answers required as input are already available. We tailor the CARE annotation platform Zyska et al. (2023) for guiding annotators through the assessment while reading the paper. Each step consists of the reasoning step’s name, prompt, detailed description, and an example, plus the prior answers or paper text serving as the input. We provide guidelines and a training session before the study (see Appendix A.2.2).
Annotation study
In the junior stage, the annotation setup was validated within a university course, where $36$ postgraduate students in the biomedical field participated voluntarily. After minor changes to the guidelines and annotator training, the senior stage involved $8$ further doctoral and postdoctoral researchers in the biomedical field, as well as $2$ postgraduate students who received dedicated training in biomedical paper assessment. After the study, we post-processed the junior data to filter out low-quality items (see Appendix A.2.2).
Dataset
The resulting dataset consists of a junior subset with $4$ annotated papers and a senior subset covering $18$ papers. At least three annotators have annotated each paper and workflow step; $11$ papers have annotations by five different annotators. The resulting dataset offers a unique opportunity for the empirical study of human decision-making and reasoning grounded in the text. The following analysis serves as a template to inspect any NL-DAR problem, showcasing the versatility of our framework. The basic statistics of the entire dataset are given in Table 1.
Dataset Statistics
# step answers $4371$
# workflow executions $93$
# papers $22$
vocabulary $7962$
# tokens $240110$
avg. # tokens per answer $60.04$
Krippendorff’s $\alpha$ $0.424$
Table 1: Dataset statistics (senior + junior subsets). Krippendorff’s $\alpha$ calculated on boolean nodes.
Boolean Decision SCM
To support the analysis, we also construct a condensed version of the workflow that consists only of the boolean variables while maintaining causal ancestry relations. Working on boolean variables makes standard causal estimation methods tractable. We approximate structural equations with Gaussian Process classifiers Rasmussen and Williams (2006). Appendix A.3.2 provides details and reports on the goodness of fit.
4.3 Analysis
As a first step towards assisting humans in NL-DAR, we aim to learn how humans perform the task. With the populated SCM, we can empirically investigate human decision-making during biomedical paper assessment. While the workflow steps are fixed during the annotation study, the answers to individual steps can and do vary. We evaluate the variation in human answers in the workflow, indicating steps that benefit from AI assistance.
Answer variability
To investigate the variability in the boolean decisions, we compute the inter-annotator agreement on the steps in the boolean decision SCM. Krippendorff’s $\alpha$ per step and across papers is $0.42$, indicating moderate agreement. This is in line with agreement levels reported for peer review ratings Bornmann (2011); Dycke et al. (2021) and typical for subjective annotation tasks. To quantify variability in the natural language answers, we follow recent studies in human answer variability Giulianelli et al. (2023) and quantify the noise induced by background factors in terms of the lexical variability (using lemma overlap), the syntactic variability (using part-of-speech bi-gram overlap), and semantic variability (using cosine similarity on SBERT embeddings Reimers and Gurevych (2019)). We compute the metrics between all annotators per step and paper and then take the average. Figure 3 shows the answer variability distribution for all steps. Notably, the relatively low lexical and syntactical similarity suggests that annotators employ their individual style and register and do not use template language, as has been observed for human-generated explanations in other studies Camburu et al. (2020). The semantic similarity, on average, lies at $0.55$, similar to other open-ended text production tasks, such as story generation Giulianelli et al. (2023). Overall, humans diagnose biomedical papers similarly, but they raise different points. The impact of background variables is reasonably low, suggesting that no important factors were left unmodelled in the SCM.
Sources of variation
We hypothesize that the answer variability changes throughout the workflow graph. Figure 4 shows the semantic similarity distribution for the different node types (infer and extract) and inference steps that explicitly invoke the annotators’ background knowledge (infer-knowledge), such as "From your experience, which plot properties are common for the given plots?". Answers to the steps based on background knowledge deviate the most, whereas extract steps show the highest similarity. The summarization-like extraction steps invoke high consistency. As infer-knowledge steps heavily depend on the individual’s knowledge, prior experience is one major background factor leading to variability. Considering variability as a proxy for step difficulty, AI support during paper assessment on background knowledge shows great potential. We provide additional analysis in Appendix A.3.1, including the propagation dynamics of variability through the causal graph.
Figure 3: Average lexical, syntactic, and semantic similarity per step and paper across all samples. Vertical dotted lines are mean values.
Figure 4: Average semantic similarity by node type. Vertical dotted lines are mean values.
Impact of decisions
The previous analysis focuses on the background variables. We now turn to the explicitly modeled variables of the SCM. We simulate $200$ samples using the boolean decision SCM trained on human data. Using average causal effect estimation (ACE) (Section 3.1), we compare the effect of interventions on different steps on the final verdict. For each node $x$ with a path to the final node $y$, we compute $\textit{ACE}(x,y)$. Table 2 lists the ACEs for selected steps. step33 (Alignment of Conclusions and Research Questions) has the highest positive impact on the final verdict. This means if the findings align well with the posed questions, the final verdict turns more positive. However, the clarity and style of writing also play an important role (step48). As the ACE directly measures the strength of causation (not an associational relation), we can conclude that annotators show a positive bias towards well-phrased papers regardless of the scientific content. This is consistent with findings in scholarly peer review Lee et al. (2013).
Counterfactual analysis
Finally, we explore the use of the SCM for counterfactual analysis. How would one need to change the underlying paper to turn the final verdict positive? We consider scenarios where an annotator gives a negative verdict ($25$ instances across $12$ papers) and study the effect of counterfactual interventions. As studying all possible $2^{18}$ interventions is intractable ($18$ steps in the boolean decision SCM), we focus on steps related to figure quality ($5$) which we call $G$. For all combinations of the figure-related interventions $C=[f_{x}=i_{x}:x\in G]$, we determine $P_{M_{C}}(t=\textit{True})$ for the final verdict $t$. We find that in $40\%$ of the scenarios, the annotator would have kept their negative verdict. For the other $60\%$, an intervention on a single step would have turned their overall assessment: the alignment between the results presented in the figures and the discussion of the results in the text (step12). In other words, the revision of figures would have frequently improved the paper. Overall, this illustrates the impact of specific paper traits on the overall assessment.
Step Prompt ACE
Considering the correct conclusions and the paper’s research questions, do the conclusions fully answer the research questions? (step33) $0.37$
Overall, do the conclusions appear relevant (i.e. in-scope of the paper and its scientific field)? (step46) $0.20$
Is this paper written clearly and concisely? (step48) $0.20$
Is this a method paper? (step4) $0.02$
Are the provided figures the common choice for the given methods? (step19) $-0.01$
Is the result reporting overall adequate for the given methods? (step20) $0.00$
Table 2: Average causal effects (ACE) of workflow steps on the final verdict on the paper. We list nodes with the highest (top) and lowest (bottom) absolute effect.
5 LLM Assistance
The NL-DAR framework allows us to decompose complex diagnostic reasoning tasks, opening new opportunities for targeted AI assistance. We illustrate this by using our data to address the following research questions: (A) Do LLMs perform biomedical paper assessments similar to humans? (B) Can LLMs act as effective AI assistants for humans?
Setup
We split the dataset into $20$ test papers ($4089$ answers) and reserve $2$ papers ($282$ answers) for prompt development. We expmeriment with LLama 3 (8B, 32k context window) AI@Meta (2024), Mixtral Jiang et al. (2024), GPT-3.5 turbo Brown et al. (2020), and GPT-4o OpenAI (2023). We design two prompt templates for each LLM, one for the extract steps and one for the infer steps, and perform the experiments zero-shot. Appendix A.4.1 lists details on the LLMs and prompts. We consider BERT-F1-Score Zhang et al. (2020) to measure the general similarity of answers, SummaC Laban et al. (2022), and TRUE Honovich et al. (2022) to measure factual alignment, and compute the F1 score for boolean decisions. We compute the automatic text similarity metrics relative to all annotators’ answers and report the averages.
LLM Diagnosis Performance
We use the SCM created in Section 4 as a scaffolding to run each LLM as a program Schlag et al. (2023); Dohan et al. (2022). We call the LLM per step and provide the answers to the parent steps as input, thereby imposing the SCM’s reasoning structure. As input to read steps, we provide the paper text and replace figures and tables with human descriptions for language-only LLMs, and provide the figures and tables as images to GPT-4o. Table 3 (top) outlines the LLM program performance. The human baseline is estimated using leave-one-out evaluation, which compares the answer of each annotator to the others and averages across all comparisons. We compute the F1 score relative to the majority label for each step as the ground truth.
The GPT models and Mixtral achieve high scores on factual consistency (SummaC and TRUE), marginally above the human baseline. However, BERT-Score ranks LLMs consistently worse. The decision-making performance on the boolean nodes lies below the average human performance by a large margin. We manually examine the discrepancy between high factuality ratings and low decision alignment in Appendix A.4.2, showing that humans reserve negative decisions for extreme cases and focus their answers on a minimal set of facts relevant to the individual assessment. LLMs, on the other hand, tend to restate inputs, split up facts into small components, and make harsher decisions. These factors might affect the factuality metrics favoring the seemingly more extensive and specific LLM answers. Overall, LLMs produce answers with high factual overlap to the average annotator but fall short in decision-making.
Simulating Human-LLM Collaboration
The NL-DAR framework allows exploring simulated human-AI collaboration scenarios. We now assume the human oversees and corrects an LLM at every step, such that the input to the LLM is the human answers from the parent steps or the human-selected paper excerpts for extract steps. We exclude steps based on figures and tables to compare all LLMs. This tests the ability of LLMs to adjust to the personal diagnosis approach of an individual human. Table 3 (bottom, ${io}$-condition) shows the results. As there is a unique answer for each input and step, we compute the F1 score relative to the respective annotator’s decision and not the majority vote as before. The human baseline is a lower-bound estimate for human performance in this scenario. [^5] For a realistic estimate, humans would need to re-annotate steps based on the input of their fellow annotators. For all LLMs, we see a notable performance increase across the board. LLMs effectively adapt to the specific perspective of an annotator as expressed by their personal inputs to each step, generating more aligned answers.
Error Propagation
The stark performance difference between the use of LLMs as a program and human-LLM joint performance suggests error propagation as a factor. To test this, we discard the graph structure entirely and call GPT-4o for each step in isolation using the prompt and the full paper as input. Table 3 shows the results on the last line. We see that the performance drastically increases compared to GPT-4o running as a program but lags behind its performance with human oversight. This strongly suggests that errors on earlier steps in the workflow deprive later steps of information or provide incorrect information for prediction. Overall, our findings show promise for LLMs as diagnostic assistants during paper assessment. More research on their integration with the SCM of an NL-DAR problem is needed. Graph-of-thought Besta et al. (2024a) and similar thought architectures Yao et al. (2023); Besta et al. (2024b) are promising to mitigate the effect of error propagation by backtracking.
BERT-F1 SummaC TRUE F1
human $\textbf{0.799}_{\pm.06}$ $-0.151_{\pm.30}$ $0.151_{\pm.27}$ 0.859
Llama3 Prg $0.752_{\pm.10}$ $-0.274_{\pm.36}$ $0.098_{\pm.30}$ $0.151$
Mixtral Prg $0.761_{\pm.09}$$\textbf{-0.149}_{\pm.26}-$ $0.120_{\pm.32}$ $0.553$
GPT3.5t Prg $0.750_{\pm.11}$ $-0.165_{\pm.33}$ $\textbf{0.159}_{\pm.37}$ $0.513$
GPT4o Prg $0.779_{\pm.07}$ $-0.196_{\pm.29}$ $0.133_{\pm.34}$ $0.696$
majorityio $0.854$
humanio∗ $0.799_{\pm.06}$ $-0.158_{\pm.29}$ $0.150_{\pm.27}$ $0.801$
Llama3io $0.786_{\pm.07}$ $-0.141_{\pm.30}$ $0.145_{\pm.35}$ $0.657$
Mixtralio $0.794_{\pm.07}$$\textbf{-0.077}_{\pm.27}-$ $0.161_{\pm.37}$ $0.822$
GPT3.5tio $\textbf{0.805}_{\pm.07}$ $-0.125_{\pm.34}$ $\textbf{0.214}_{\pm.41}$ $0.789$
GPT4oio $0.795_{\pm.07}$ $-0.094_{\pm.28}$ $0.194_{\pm.40}$ 0.876
GPT4o§ $0.775_{\pm.07}$ $-0.156_{\pm.28}$ $0.187_{\pm.39}$ $0.842$
Table 3: Performance under varying conditions. $\cdot^{\dagger}=$ LLMs run as a program. $\cdot^{io}=$ LLMs receive the human input for each step. GPT4o§ receives prompts and the full paper text without the graph structure. We average over eligible steps and papers, with the standard deviation in subscript. For all metrics, the higher, the better.
6 Conclusion
We introduced NL-DAR as an important class of abductive reasoning problems and proposed a framework rooted in causality theory to approach NL-DAR problems. We outlined a practical guideline and toolkit to define and analyze NL-DAR problems. Our novel large-scale human dataset instantiates the framework for biomedical paper assessment. Our experiments showcase the versatility of our framework for analysis and LLM assistance. Our insights pave the path towards a broad study of natural language diagnosis tasks on the interface of causality, human-AI collaboration, and LLMs.
Limitations
Framework
Our framework for modeling NL-DAR problems is designed to be general and applicable to any natural language diagnostic reasoning task. We demonstrate its application and share the insights from our case study on biomedical paper assessment. More research is needed to validate the application of the NL-DAR framework for other diagnostic tasks.
At the core of our framework lies the definition of NL-DAR problems via a structural causal model. Here, we assume that the final verdicts arise when consuming the input text and do not exist on their own. This perspective is subject to philosophical discussion and most prominently advocated by Roland Barthes Barthes (1977). In practice, this means we assume that a text is consumed, causing the observed verdict (e.g., "the paper is a valuable contribution.") and that the reasoning steps in between give rise to this verdict. However, one might argue that there should be an objective verdict; for example, for paper assessment, a paper is either objectively correct or not. Likewise, one could argue that the verdict comes first, and the diagnosis is the post-hoc justification for it. While we acknowledge these points, we emphasize that the practical application of our framework in the case study supports our design choices in that diagnostic problems in natural language give rise to variance in the final verdict and in that the expert interviews support the directionality of causation that underlies our framework.
Case Study
The case study of this paper focuses on the biomedical domain. We believe that the developed SCM will be generally relevant to paper assessment in other scientific domains and can be used as a starting point for domain adaptation through expert interaction. Yet, the findings on the factors that shape human diagnosis are limited to the biomedical domain; researchers from other fields might put a different focus during the assessment; for example, the contribution of datasets and open resources is not considered in the biomedical paper assessment workflow, but is an important criterion in the assessment of NLP papers.
The data collection aims to cover a wide range of biomedical papers from various institutions and world regions to avoid biases toward a particular style of writing, scientific agenda, or scientific sub-community. However, some topical bias is inevitable to ensure relevance to the annotators’ expertise. We ensured high redundancy during annotation to avoid biasing the dataset to the idiosyncracies of individual annotators. Yet, the given sample of papers and the annotator pool might still introduce a certain level of bias to the dataset.
In the evaluation of LLMs as assistants for diagnostic reasoning, we rely on a range of automatic metrics for text generation. As our findings suggest, these metrics might not cover all aspects of a suitable answer in the workflow; for instance, they seem to favor more extensive answers, putting efficient but concise responses at a disadvantage. Likewise, the human baseline is a conservative estimate of true human performance; it is important to interpret our findings accordingly. Automatic metrics that consider the reasoning context of an answer, similar to the evaluation of chain-of-thought reasoning in LLMs Hao et al. (2024), are a promising direction of research to address these issues.
Ethics
This study has been approved by the university’s ethical board under the proposal number EK 41/2024. All senior annotators and experts participating in the study received adequate compensation within their regular employment in the research project. For the junior annotators, the workflow answers were collected as part of a university course aiming to teach students critical reading skills. The participation in the annotation study did not impact the students’ grading in the underlying course. All annotators could choose freely to reject participation in the study, and all human data was collected based on explicit consent.
We explore the use of LLMs to automatically diagnose and evaluate biomedical papers. This technology has potential dual uses in peer review. A malicious party could use the LLM to skip the manual paper assessment process, which arguably requires the most effort during peer review. This is clearly problematic because LLMs cannot entirely substitute a human researcher, leading to incorrect or biased peer reviews. We stress that the output of such a system would be the answers to the reasoning steps and not a full peer-review report, which also covers questions to the authors or suggestions for improvement. We argue that the risk of misuse of this technology is relatively low. At the same time, the benefit of enabling new research on the reasoning underlying peer review and using paper assessment workflows as a teaching or assistance device for reviewers is considerable, and we believe that the benefits of the proposed technology outweigh the risks.
Underlying our dataset is a set of publicly available papers with open licenses (CC-BY, CC0, …) that allow the redistribution and extension of the contents. However, as we publish subjective assessments of the papers by our study participants, this could potentially harm the reputation of the publication authors when they are assessed unfairly. We highlight that this is not the intended use of our dataset, which is exclusively meant for research on natural language diagnostic reasoning. We partially mitigate this risk by having at least three different researchers make annotations per paper: as a result, the collected dataset shows an opinion spectrum rather than a singular verdict on the paper. Hereby, we deem the risk of harming authors to lie below the day-to-day risk for reputational damage in regular scientific publishing; peer reviews are commonly released to the public, and papers are discussed (and criticized) publicly, for instance, on social media and in subsequent publications.
Appendix A Appendix
A.1 Constructing an SCM for Biomedical Paper Assessment
A.1.1 Interviews and Formative Study
We report on additional details of the interviews and the formative study used to elicit the workflow from the domain experts. The process was divided into four major steps: before the interviews, the actual interviews, consolidation, and the formative study. For all stages, we worked with two biomedical experts on post-doctoral and professorial levels, respectively. Both have reviewed numerous publications during their careers and have extensive experience in the field.
Before the interviews
Before the interviews for deriving an SCM, we determine the critical points during text assessment and characteristics of the paper genre to ground all further steps on this information. After interrogating the two experts in a joint session we converge to the following insights on the domain at hand.
First, there are two general types of papers in biomedicine relevant to their work: natural science papers, which investigate natural phenomena, and method papers, which propose a new approach to conducting experiments. Second, biomedical papers follow a relatively rigid surface structure with an introduction including background information, the methods section as a detailed experimental log, the results section reporting the exact measurements, and the discussion section followed by an overall conclusion. The results and discussion section can be inter-mingled. Third, figures and tables play a major role in the understanding and assessment of a paper. Fourth, methodological soundness and consistency of results are key indicators during text assessment. Finally, several heuristics exist to detect bad experiment design, such as the lack of controls or suspiciously high standard deviations on measurements.
Interviews
We conducted the interviews with each expert in isolation and orienting towards best practices of interview design by Knott et al. (2022). To structure the interview, we discussed the assessment criteria and process based on reviewing checklists of the field, starting with a general and intuitive description of the assessment process and increasing the level of detail. For the first interview, we considered the manuscript checklist by Seals and Tanaka (2000) conceptualized for physiology research. For the second interview, we use the Portland Press review checklist [^6] https://portlandpress.com/pages/peer_review_checklist (accessed 08/2024) for biomedical sciences. We opted for two different checklists to encourage different perspectives during the workflow elicitation and mitigate the impact of the interviewer’s biases.
The result of these interviews is two different step-wise manuals on how to iterate through a biomedical paper and assess its methodological soundness. We identify the there types of actions taken during assessment (read, extract, infer) which align between both manuals. While some of the subordinate steps deviate and the researchers put a different emphasize on each of these aspects, the overall assessment approach is very similar. In summary, first, the whole article is skimmed to get an understanding of its contents. Then, the experts check the contextualization with related work and background information in the introduction. Afterward, they skip to the results section, which includes figures and tables. They check the alignment of results and whether the reporting seems plausible and follows best scientific practice. Subsequently, they read the discussion in detail and analyze if the provided interpretations and conclusions align with their own readings of the results and their context in related work. Then, they verify that the method section covers all important methods and that they are described correctly. Finally, they check the conclusion section, abstract and title to verify that they align with findings and methodology described in the other sections. Both experts agree on the temporal ordering of these major steps during assessment.
Consolidation
After the interviews, we consolidate the sequential manuals into one manual covering all the mentioned points by both researchers. We then interpret the temporal ordering and the sub-ordination relationship between major and minor steps as causal dependencies to answer follow-up questions. For instance, to answer a sub-ordinate step, "Is the description of the state-of-the-art complete?" we require the input from the super-ordinate step, "How do the authors describe the state-of-the-art in the field?".
This results in a workflow of $106$ reasoning steps covering all sections of the paper. We discuss the dependency structure and overall workflow with one of the domain experts and incorporate suggested changes. We use the resulting workflow as input for the formative study.
Formative Study
We perform the formative study with three PhD students from the same research lab as the two senior experts. Following a similar setup as described in the final annotation study (in Section 4.2), we linearized the workflow. We asked them to assess a paper selected by their preference, providing answers for each step. For the formative study, the annotators work without any specific annotation interface and simply append the answers to a text document. We summarize our findings from this formative study in the following.
First, the annotators did not finish the full workflow of $106$ steps within the three hours of the study, suggesting that the workflow should be subselected to keep the annotation practical. Second, the annotators tended to ignore the causal dependencies between the steps, meaning that they answered several steps in isolation based on the paper and not their own prior answers. The main obstacle for this lied in the lack of a tailored annotation interface that presents prior answers and enforces the causal structure. Third, annotators need training to get an overview of the workflow graph to avoid redundantly stating similar facts across multiple steps. An easy option to revise prior answers might also be interesting for that purpose.
In response to these observations, we designed the final evaluation study setup to collect the dataset underlying the SCM (as described in 4.2). More details on the annotation interface can be found in the Appendix A.2.2.
A.1.2 SCM for Biomedical Paper Assessment
We give a more extensive overview of the graphical structure, the specific steps, and their prompts in the biomedical paper assessment workflow.
Workflow Steps
The full causal graph of the biomedical workflow is illustrated in Figure 5. Tables 4 to 6 list all workflow steps, including their type and prompt. The description and example are omitted for readability. We defer the reader to the associated codebase for the complete workflow step specifications.
Graph Structure
The graph, as shown in Figure 5, is a directed and acyclic graph. The average degree per node lies at $3.57$; the mean number of parents lies at $1.78$; there are $3$ roots in the graph (the read) nodes; there are $4$ terminal nodes of which one is the final verdict, whereas the others are the assessment of clarity and the high-level summary based on the skim. These only exist to ease the annotation and have no meaning for the overall SCM modeling of the DAR problem.
Upon qualitative inspection of the graph, it becomes clear that the two major sections (results and discussion) form a dense network of descendants, while there are only a few connections between the nodes belonging to the different sections. Only the nodes closer to the final verdict share parents belonging to different sections. This allows the graph to be divided into $5$ layers of nodes, where each layer contains only nodes that can be assessed in parallel according to topological ordering. We make use of this property for the batch-wise computation of the LLM programs during our experiments described in Section 5.
Boolean Workflow Version
For our causal analysis of impact factors on decision making (see Section 4.3), we convert the full text-based workflow into one that consists only of the boolean nodes to estimate the structural equations from data.
Figure 6 shows the causal graph of the resulting boolean workflow. This graph maintains the causal ancestry relationships of the original graph. More formally, the output graph maintains the transitive closure of the original graph on the subset of selected nodes. To arrive at this, we substitute all text-only nodes (such as extract nodes) with an edge, which means that we basically "skip" this step but maintain the causal relationship from their parents to their children. For nodes that would be substituted and lie at the end of a path, we drop the nodes. This is the case for read nodes. However, as all read nodes depend on the paper, we add an additional joint dependency of all children of read nodes to the node assessing the clarity of the paper. We make the assumption that the read nodes of different sections are independent except for the clarity of the paper acting as a confounder.
Figure 5: Representation of the full workflow. For each step, the prompt is provided. The color of each node indicates the node type. read nodes are red, extract nodes are green and infer nodes are blue. The turquoise nodes are infer nodes that rely on background information not present in the paper. The edges are colored to match the parent node type and increase the readability.
Figure 6: Representation of the boolean workflow. For each step, the name and ID are provided.
Id Name Type Prompt
step001.42 Skim read Skim the whole paper briefly (10-15min).
step002.44 High-level Idea extract Summarize the high-level idea of the article in 4-5 sentences.
step034.50 Research Questions extract What research questions (a.k.a. hypotheses) do the authors investigate? Please highlight and list
step003.43 High-level Methods extract Summarize the used methods of the article briefly and on a high-level.
step004.8 Method Paper Type extract Is this a method paper (as opposed to an original research paper)?
step048.x Overall Clarity of the Paper extract Is this paper written clearly and concisely?
step006.4 Results Section read Read the Results section (and associated figures in the appendix) in detail. Highlight the caption of the section to indicate where it starts.
step007.7 Figures extract Highlight all figure captions relevant to the result section and summarize the figures briefly without interpreting them. If there are no figures, respond with ’none’ as a list step.
step016.25 Expected Figures for Methods infer What types of plots/figures/tables would you expect for the methods used in the paper?
step019.60 Choice of Figures infer knowledge Are the provided figures the common choice for the given methods?
step017.58 Missing Figures infer Are there any important figures missing?
step008.6 Results in Figures extract Consider all figures/plots relevant to the results section. Summarize the results implied by the figures; i.e. specific observations on the plotted data.
step009.71 Typical Figure Properties infer knowledge From your experience, which plot properties (such as standard deviation, value ranges, …) are common/expected in your research field for the given result plots?
step010.72 Plausibility of Figures infer Considering how you imagine typical plots and the actual plots you observed in the results section, do the plots seem plausible?
step011.28 Results in the Text extract What results/observations do the authors report in the TEXT based on the experimental data and figures? Please highlight and list.
step012.41 Matching Figures and Text infer Considering what the authors report in the text and the results you read from the figures, do they match?
Table 4: The first steps of the workflow with ID, type, and prompt.
Id Name Type Prompt
step013.29 Summary of Results infer Based on your prior observations regarding the results the authors report, the figures imply and their consistency, please summarize the main results in a short paragraph.
step014.5 Figure Tempering extract Do the figures show signs of tempering?
step015.73 Truthful Figures infer Do the provided figures appear truthful and adhere to best scientific practice?
step018.59 Result Reporting Completeness infer Is the result reporting complete (i.e. no important results are ommitted)?
step020.74 Overall Impression Result Reporting infer Is the result reporting overall adequate for the given methods?
step021.12 Discussion Section read Read the discussion section in detail. Highlight the section caption to indicate where it starts. If there are also interpretations and conclusions in the result section, please consider that too.
step022.20 Interpretations extract Which interpretations do the authors provide? Interpretations = explanations for a specific experimental result
step023.19 Direct conclusions extract Which direct conclusions do the authors draw? Direct conclusions = general knowledge claims based on an experiment
step024.18 Projective conclusions extract Which projective/speculative conclusions do the authors make? Speculative conclusions = opinions = plausible claims that are not fully backed by the data
step025.26 Method confounders infer knowledge Which confounding factors for the used methods are known to you?
step026.30 INVALID interpretations infer Which interpretations of the authors are invalid considering potential confounders of the experiments and actual reported results?
step027.33 VALID interpretations infer Which interpretations of the authors are justified by the actually reported results?
step028.34 ALTERNATIVE interpretations infer knowledge Considering the valid interpretations, can you imagine additional or alternative interpretations that are equally justified by the data?
step029.98 Incompleteness of interpretations infer Considering your alternative and additional interpretations, are they actually missing in the valid interpretations of the authors? Judge the overall completeness of interpretations based on this.
step030.35 VALID conclusions infer Considering the authors’ direct conclusions, which are valid and derive validly from the correct interpretations?
step031.31 INVALID conclusions infer Considering the authors’ conclusions, which are invalid given the interpretations?
Table 5: The intermediate steps of the workflow with ID, type, and prompt.
Id Name Type Prompt
step033.99 Completeness of conclusions infer Considering the correct conclusions and the paper’s research questions, do the conclusions fully answer the research questions?
step035.17 Related work results extract What related work RESULTS do the authors report? Please highlight in the paper and summarize the reporting briefly.
step36.14 Related work results differences extract Do the authors observe and explain differences and commonalities of their results compared to related work? Please highlight and summarize briefly.
step37.102 Related work results missing infer knowledge Based on your knowledge of the related work and the related work results the authors report, is there anything missing? In that case, please list the results of related work that are missing.
step38.103 Results contextualization infer Are the results appropriately contextualized in the related work? Please also consider the paper type to weight your judgement.
step39.104 Results plausibility infer Are the experimental results plausible given the results of related work? Please weight your judgement by the paper type.
step44.105 Result validity infer Overall, do the RESULTS appear scientifically sound? Consider the adequacy of reporting, plausibility of results, and the related work context.
step40.13 Related conclusions discussion extract Do the authors discuss and explain differences and commonalities to CONCLUSIONS in related work? Please highlight in the paper and summarize the discussion for each aspect briefly.
step42.x Related conclusions comparison adequacy infer Is the discussion of the related work CONCLUSIONS appropriate? Consider if there are alternative explanations possible or if there are errors in the argumentation.
step41.15 Related conclusions missing infer knowledge Based on your knowledge of the related work and the related work CONCLUSIONS the authors discuss, is there anything missing?
step43.96 Conclusions contextualization adequacy infer Overall, are the conclusions properly contextualized with related work? Take the paper type into account to weight your response.
step45.106 Conclusion validity infer Overall, do the correct conclusions seem scientifically sound and answer the research questions? Consider if conclusions cover all research questions, the correctness of underlying results, and the authors’ projective conclusions.
step46.100 Conclusion relevance infer Overall, do the conclusions appear relevant (i.e. in-scope of the paper and its scientific field)? Consider if the conclusions answer the research questions and the related work context.
step47.x Overall impression infer Overall, do you consider the paper a valuable contribution to the scientific community?
Table 6: The final steps of the workflow with ID, type, and prompt.
A.2 Data Collection
A.2.1 Underlying Paper Dataset
We crawl the underlying papers from bioRxiv and ensure a high sample diversity along relevant dimensions. Here, we outline the details of this process.
Data Acquisition
We retrieve the index of all papers on bioRxiv using a custom crawler based on the recent paper list [^7] https://www.biorxiv.org/content/early/recent, accessed 08/2024. We deduplicated the list, resulting in roughly $100$k in preprint URLs. We then retrieve all paper meta-data, including the title, authors, corresponding author institution, type of study, version, category tag, publication information, license, abstract, and the number of versions of each article, through the bioRxiv API. We augment each entry by the country and world region of the authoring institution [^8] We match the institutions via the publicly available university domain mapping https://github.com/Hipo/university-domains-list and the paper length estimated from the number of PDF pages.
[…]
Sample Diversity
Based on this dataset, we aim to select a representative subset along the dimensions of the world region of the authoring institution, the number of versions that exist of the preprint, and the publication status. The latter two criteria should ensure that there are preprints of varying quality in the samples, whereas the first should ensure that the papers come from diverse institutions with different publication records and writing styles. For this purpose, we take a stratified sample of $200$, which is then manually inspected and filtered by one expert annotator, resulting in $57$ eligible papers, of which $22$ were annotated. Figures 7 to 9 show the sample diversity of the final dataset along the key dimensions as compared to the initial stratified sample of $200$ papers. While observing a light skew towards accepted papers, all world regions, versions, and publication statuses are present in similar proportions as in the full sample. All papers are written in English which is the lingua franca across science.
Figure 7: The ratio of each world region in the full paper dataset and the subsample of $22$ papers included in the annotation study.
Figure 8: The ratio of versions in the full paper dataset and the subsample of $22$ papers included in the annotation study.
Figure 9: The ratio of publication status in the full paper dataset and the subsample of $22$ papers included in the annotation study.
Data Format
For all resulting papers, we retrieve the PDF and the XML from bioRxiv. To complement the XML with figures and tables, which are represented as images, we reverse engineer their URLs and retrieve those as JPG files. We parse the XML together with the images and store them in Intertext Graph Format (ITG) [^10] https://github.com/UKPLab/intertext-graph/tree/main.
A.2.2 Annotation Study Details
Overview
We employed the following guidelines and annotation interface to ensure maximal consistency among annotators and to encourage step-wise causal thinking. As another key quality measure, we repeatedly engaged with individual annotators and the whole annotator group to calibrate and provide feedback on their annotations.
Guidelines
The guidelines consisted of the following text extended by a sequence of images explaining the annotation interface in detail.
A “reviewing workflow” is a formal way of reasoning about the quality of a paper. A workflow consists of a sequence of questions that build on top of each other. The workflow was developed with experts of the field and while it might not resemble your personal internal mindmap during peer review, it is a good way of discussing an article’s quality.
Some of the questions in the workflow simply extract information from the article or your background knowledge, others infer or combine quality judgements based on earlier answers. you will encounter the following questions/tasks in a workflow: read – here you read a section or part of the paper. extract – based on the section you read (in an earlier question) you try to extract certain information from that section. Usually this means you will need to highlight some information in the paper and write a summary. infer – based on previous answers to questions, you make a judgment on the question (usually a yes/no question) by combining the previous answers in the right way.
Your task in the annotation study is to strictly follow the reviewing workflow as you are guided through its questions step-by-step. While this seems unnatural at first, this makes sure that all participants follow the same reasoning, while still answering each question naturally and from their subjective point of view. During the annotation there are several things to keep in mind, which are explained in more detail below. Most importantly, you should try to annotate each question based on precisely the input that is provided to you during the workflow; do not answer questions in isolation and by a vague general feeling of the paper.
Reviewing Workflow Example
In the following you will be guided through a dummy workflow to make you more familiar with the annotation interface and the concept of workflows. Before you dive into CARE, please have a look at the right example for the underlying logic of workflow. When you do the annotation and in this mini-tutorial this graph of dependent questions is converted into a questionnaire.
Please click on the following link to go through the tutorial. If you should have problems with stepping through the workflow, please refer to the section Annotation Interface or talk to the study coordinators. We recommend Firefox for using CARE.
Terminology
Some terminology re-occurs in many questions of the workflow and there is some room for misinterpretation. Please briefly consider the following terms and “our” definition for them.
  • Research question (aka hypothesis): A research question or hypothesis = a claim (possibly phrased as a question) on the system under investigation which the authors try to verify or falsify; example: ’is brain cancer caused by this DNA sequence?’
  • Method (aka approach): The methods or approach of a paper = the (standardized) tools, technologies, and actions the authors undertake to approach the research question; example: modular cloning, Gibson assembly, …
  • Result: The results of a paper = the data measured/generated during the experiments of the paper. The results comprise more or less the raw measurements, possibly aggregated by statistics, but not interpreted.
  • Interpretation: An interpretation = based on results (raw data) the authors provide an interpretation, i.e. an explanation or conclusion within just the scope of this specific experiment run, but without claiming correctness of the statement for the whole natural science world; example: the experiments show an upward trend in variable X.
  • Conclusion (aka finding): A conclusion = based on the interpretations of experimental results and the validity of the experimental setup, the authors generalize a claim from the specific experiment to the natural science world; example: variable X and variable Y correlate.
  • Method Paper Type: A method paper introduces a new experimental setup / technology / tool / technique to solve a practical research problem; e.g. discovering a new RNA lever to enable follow-up research or designing an algorithm to measure quantity X. The central research question usually is “Can we develop a new tool for task X that has properties Y?”. The method of such a paper comprises the newly proposed tool/setup/technique. The results cover the measurements on desired properties Y on task X (e.g. accuracy) of the tool. The interpretations analyze the significance of these properties (compared to related work) and the conclusions are usually claims about whether the new tool is an improvement to existing tools along certain dimensions and for which scientific endeavors it should be applied.
  • Original Research Paper Type: An original research paper investigates a research question on natural science phenomena; e.g. checking if two molecules interact with each other or searching for an explanation for certain behavior in cells. The central research question can be diverse; they are usually a causal question (“does A cause B?” / “why does X behave as observed?” / etc.). The method of such a paper comprises the experimental setup based on any prior tools/approaches and how they are combined. The results cover measurements on the observed variables and the interpretations analyze these results to form specific claims about the experiments. The conclusion usually puts the interpretations into perspective and makes claims about the underlying natural science phenomenon.
Principles
  • English Language – please answer the questions only in English.
  • Fluency – In general, answer all questions with fluent text.
  • Explanations – For yes/no answers, please always first provide your assessment (yes / no), but then also add an explanation that describes how you arrive at this answer and contextualizes/constraints your assessment. This is especially important when you cannot give a clear yes/no response. E.g.: Q: “Are there outdated references?” (good) A: "Yes. There are some references to methods that are clearly outdated: they are more than 5 years old. However, the references to the most relevant work on the same issue are recent, so I judge the references to be overall sufficiently new.” (bad, trivial explanation) A: “Yes. Because the references are not outdated.”
  • Good reasoning (what-if) – on infer questions, where you combine and weigh prior answers to respond to the new question, you should ask yourself, what input answers would turn my judgment positive/negative. Likewise, you should check that all information in your answer really comes only from the inputs.
  • Uncertain Flag – You can mark an answer as uncertain (flag icon), if you really don’t feel able to answer the question. However, use this option very sparingly!
  • Highlighting – For the extractive questions you should try to highlight whole sentences and not on sub-sentence level. Ideally, for each aspect you extract you can narrow down the highlight to at most 4 consecutive sentences – avoid selecting whole paragraphs, let alone pages. If you need to highlight across two pages (not possible in CARE), please simply make two highlights.
  • Reading Tasks – if a reading task points you to a certain paper section, do consider if this section is present as such or if you should peek into others to answer the questions; e.g. some discussion might appear in results section.
  • Comprehensive Answers – Remember that all your answers to questions will form the input to later steps in the workflow. This means your answers should be comprehensive to allow all follow-up judgements. If you notice at a later question that your earlier answer is insufficient, do not hesitate to jump back to that earlier question and flesh it out. With some practice in using the workflow, this will become less frequent.
  • Don’t be too picky and don’t be too nice. The purpose of the reviewing workflow is to perform a mostly neutral assessment of the paper at hand. It is not about “destroying” the paper (even if you disagree with the goals/ideas/… of it) and you shouldn’t naively trust everything the authors say or get distracted by the authors’ good or bad writing style.
Examples
Q: What is the pool of related work you know?
Bad: I somewhat remember there are some other works that deal with that phenomenon. → informal, vague
Better: There is a list of related works that deals with the same phenomenon: X, Y, Z. → formal, precise
Better: I am not familiar with the detailed related work of the field, however works X, Y, Z appear similar to the scope of this work. → formal, objective
Q: Are there references missing? (based on 2 and 3)
Bad: Yes. → no explanation, no context
Bad: No, it really sounds like they cover all bases. → informal, not derived from 2 + 3, but a general feeling on the paper directly
Better: Yes. While they reference the most important works (X, Y), they omit Z, which also investigates <subject>. → contextualized, explained
Q: Are there outdated references? (based on x, y)
Bad: Yes → no explanation, no context
Bad: Yes, because there are outdated references → trivial explanation
Better: Yes. The references to the methodology are all older than 10 years. In the given fast-paced area of research such references are generally outdated. → contextualized, explained
Annotation Interface
Screenshots of the annotation interface are shown in Figures 10 and 11. On the right, the annotators can control how they step through the workflow. They can jump back to earlier steps if needed. In the view for infer items, annotators cannot see the paper and have to base their answer on their prior answers.
Figure 10: The annotation interface for an extract step. The paper with highlights is on the left; on the right, the user receives the task for the given step. They can answer in a structure or free-text form, depending on the question. Annotators can skip through the steps, flag them, and jump to prior steps as needed.
Figure 11: The annotation interface for inference steps. Here, a modal opens above the paper to block the view and encourage annotators to concentrate only on the provided inputs of prior steps (on the left) to write their answers.
Annotator Demographics
We report selected annotator statistics including demographics for reproducibility. Due to the relatively small pool of annotators, we report the annotators of the Junior and Senior study jointly and omit detailed characteristics that could violate their anonymity. The identities of all annotators are kept secret and we assign random pseudonyms in the dataset to avoid deanonymization.
The distribution of genders is nearly balanced between male and female; diverse was never self-reported. Nearly all annotators are from the age group $20-29$. Figure 12 reports the number of years each annotator is engaged in biomedical research in any form. Figure 13 shows the distribution of biomedical papers read per week in the whole pool. Several of the Junior annotators do not regularly read biomedical papers, but several consume them occasionally. The Senior annotators all consume papers regularly. Nearly all annotators speak English as a second language.
Figure 12: Self-reported number of years concerned with biomedical studies.
Figure 13: Self-reported number of papers read per week on average.
Dataset Post-processing
We pseudoynmize all data and bring the annotations into a unified format. Because the Junior annotators have less experience with respect to academic review and publishing, we filter this dataset by a set of quality heuristics to ensure high quality. As a proxy for quality, we manually inspect the answers of each annotator. We discard annotators’ answers that provide a drastically different summary of the article compared to other annotators and the abstract. We further discard annotations by annotators that provide consistently shorter answers than the other annotators. Finally, we discard annotations which contain multiple trivial explanations ("A is because A is.") or provide non-English answers. Overall, we filter highly restrictively to ensure high quality of the dataset. We discard annotations by $20$ of the Junior annotators.
A.3 Data Insights
A.3.1 Detailed Analysis of Answer Variability
We complement the analysis of answer variability in the main paper with a detailed analysis of variability throughout the graph.
Does variability depend on the areas of the paper that are analyzed?
When considering the semantic variability along the workflow graph, we see a weak negative correlation (Kendall’s $\tau=-0.36$) between the position of the step in the linearized workflow and the average semantic similarity. Here, the number of preceding steps seems to play a role (Kendall’s $\tau=-0.15$), but also the scope of the steps focusing either on high-level information (mean $0.60$ cosine similarity), results ($0.63$), or interpretations and conclusions ($0.48$) in the paper. This is plausible given that the discussion section of a paper gives rise to more subjective assessments than the raw results. All in all, the variability depends on the elements of the paper that are discussed.
Does variability propagate through the workflow graph?
Considering the relatively high semantic alignment between annotators on extraction tasks (mean semantic similarity of $0.60$), a lot of the answer variability has to be the consequence of deviating opinions during intermediate reasoning steps despite relying on the same evidence. To discover the impact of deviating opinions, we investigate the propagation effect of answer variability along paths of the workflow graph and for individual annotators.
We focus on answers that are substantially different from the ones of the other annotators for a step. We define an answer as substantially disagreeing (dis) if the mean semantic similarity to the other annotators lies one standard deviation below the average similarity across all papers for that step. This means, if an annotator shows exceptionally high dissimilarity with the other annotators we flag this answer as a substantial disagreement. Under this assumption, roughly $21\%$ of all answers show substantial disagreements. We can then compute the probability of propagating disagreement as $P(\textit{dis}|\geq 1\textit{ parent with dis})=0.72$. As expected, substantial disagreement propagates to the subsequent questions with a $72\%$ chance as compared to random disagreement with only $21\%$.
A.3.2 Structural Equation Approximation from Data
We report on the boolean decision SCM for analysis. Specifically, we outline the goodness of fit for the resulting SCM trained on the human-generated dataset.
Boolean SCM Representation
The model is represented as an invertible structural causal model within the do-why library [^11] https://www.pywhy.org/dowhy/v0.11.1/. The root nodes of the graph are estimated as empirical distributions, and the other nodes are assumed to be sampled with additive noise. To learn the mechanism of a node, we use Gaussian process classifiers. These assumptions allow us to estimate the structural equations from data and determine the noise variables for counterfactual queries.
Goodness of Fit
We rely on the evaluate_causal_model function of the do-why library to estimate how well the learned SCM approximates the observed human decisions to measure its validity.
The graph structure of the SCM aligns well with the data, as there are no other permutations of the DAG inside its Markov equivalence class. The average KL divergence between the distribution implied by the learned structural equations lies at $0.18$ which shows an overall close fit to the data. The invertibility assumption of the nodes is not rejected for any of the nodes except the root node step4, which determines if the paper is of the type method paper. Notably, the Continuous Ranked Probability Score is low for all nodes (below $0.31$) except for step4 at $0.39$. While this hints at the fact that the model might not perfectly reflect this factor, the overall model fits the observed data well.
A.3.3 Implementation Details
We report on the implementation details for the analysis of the SCM.
To determine the semantic similarity we use SBERT Reimers and Gurevych (2019) provided by the sentence-transformers library (version 2.7.0). We use the embedding model "sentence-transformers/all-mpnet-base-v2" trained on MPNET. For the similarity by part-of-speech tags and lemmas, we rely on spacy’s [^12] https://spacy.io/ spacy-v3.6.1 and scispacy-v0.5.3 language processing module specialized for the scientific domain en_core_sci_lg.
A.4 LLM Assistance Experiments
We report on hyperparameters and all details used for the different experiments. In total, the experiments with commercial models invoked roughly $200$ USD of API inference costs. The open models were run on three A100 GPUs.
A.4.1 Experimental Configurations
We report on the detailed LLM versions, hyperparameters, and prompts for the three experimental scenarios (automation, human-LLM simulation, and ablations on error propagation). Additionally, we report on the evaluation metric versions used.
Open LLM Configurations
We use the langchain [^13] https://www.langchain.com/, langchain-v0.2.8, langchain-openai-v0.1.19 and huggingface transformers [^14] https://huggingface.co/, transformers-v4.35.2 libraries to load the open LLMs for inference.
We use the instruction-tuned LLama-3 model with 8B parameters and the extend 32k context window [^15] https://huggingface.co/NurtureAI/Meta-Llama-3-8B-Instruct-32k to fit full papers into the input for extract nodes. For Mixtral, we use the mixture of 8 experts with 8B parameters instruction-tuned [^16] https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1. We also tested other models including Llama-3 with 70B parameters but the context size was insufficient to fit the relevant inputs.
We run all open models with the generation configuration specified in Table 7. For Llama we add the necessary special tokens for termination.
parameter value
max new tokens 2048
temperature 0.0001
top p 0.95
do sample True
repetition penalty 1.15
Table 7: LLM Generation Configuration
Closed LLM Configurations
For the closed LLMs we use the AZURE OpenAI API [^17] https://azure.microsoft.com/en-us/products/ai-services/openai-service. We use the models GPT-3.5-turbo with a 16k context window  [^18] https://platform.openai.com/docs/models/gpt-3-5-turbo, gpt-35-turbo-0613-16k and GPT-4o [^19] https://platform.openai.com/docs/models/gpt-4o, gpt-4o with default generation parameters.
Prompts and Prompt Selection
We design two prompts for each model, one for the extract steps and one for the infer steps. Table 10 shows the prompts for LLama-3, Table 11 for Mixtral, Table 8 for GPT-3.5-turbo, and Table 9 for GPT-4o. The template parameters are substituted before inference: the parent parameter is replaced by the paper text or prior answers in the graph; the task parameter is replaced by the prompt for that step in the workflow; the description is replaced by the detailed task description of that step in the workflow; the format specifies a json object matching the expected output type (e.g. including a boolean field for boolean questions); example is substituted by the same example that annotators saw during annotation but in a suitable json format.
We first tune the prompts by manual inspection until the output format and answers appear plausible. We then tune them manually for an optimal BERT-Score on the development set of the data. We do not perform an extensive search for the best prompt templates; conceptually, we could tune a prompt for each step individually. This is infeasible within the scope of this work.
Type Prompt
extract "system": "You are a biomedical researcher assessing the quality of a research article." "human": "This is the research article text: {parents} Based on the article text, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}". "
infer "system": "You are a biomedical researcher assessing the quality of a research article." "human": "You already gathered the following insights on the paper by answering a sequence of prior questions: Your insights: {parents} Based on these prior insights, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}". "
Table 8: GPT-3.5-turbo prompts in chat format. {…} mark template parameters.
Type Prompt
extract "system": "You are a biomedical researcher assessing the quality of a research article." "human": "This is the research article text: {parents} Based on the article text and figure descriptions, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}". " "human": {img_caption} "human": {img_base64}
infer "system": "You are a biomedical researcher assessing the quality of a research article." "human": "You already gathered the following insights on the paper by answering a sequence of prior questions: Your insights: {parents} Based on these prior insights, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}". "
Table 9: GPT-4o prompts in chat format. {…} mark template parameters. The infer prompt is identical to the one of GPT-3.5-turbo. For extract steps, the image caption and data messages are repeated for each figure and table relevant to multimodal steps.
Type Prompt
extract "system": "You are a biomedical researcher assessing the quality of a research article." "human": "The research article text: {parents} Based on the article text, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}"."
infer "system": "You are a biomedical researcher assessing the quality of a research article." "user":"You already gathered the following insights on the paper by answering a sequence of prior questions: Your insights: {parents} Based on these prior insights, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}"."
Table 10: Llama3 prompts in chat format. {…} mark template parameters.
Type Prompt
extract "human": "You are a biomedical researcher assessing the quality of a research article. The research article text: {parents} Based on the article text, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}"."
infer "human": "You are a biomedical researcher assessing the quality of a research article. You already gathered the following insights on the paper by answering a sequence of prior questions: Your insights: {parents} Based on these prior insights, answer the following question: {task} More specifically, this means: {description} Keep your answer concise. Return the answer as a JSON object in the following format: {format} Here’s an example output for the given question: "{example}"."
Table 11: Mixtral prompts in chat format. {…} mark template parameters.
Evaluation Metrics
For evaluation, we use the BERT-F1-Score from the huggingface evaluate library [^20] https://huggingface.co/docs/evaluate/en/index evaluate-v0.4.2 and choose distilbert-base-uncased as the base model. We use the huggingface version of the TRUE score [^21] https://huggingface.co/google/t5_xxl_true_nli_mixture and choose google/t5_xxl_true_nli_mixture as the model. We parse the first character as the output of the model and assume a default score of zero if the answer cannot be parsed. For SummaC we use the author provided version [^22] https://github.com/tingofurro/summac summac-v0.0.4.
A.4.2 Extended Analysis
In this subsection we report on further analysis conducted to gain insights into the reasons for the measured performance of LLMs.
Decision Making
Humans mostly lean towards positive decisions ($67\%$ ’yes’ answers), whereas most LLMs are more restrictive ($44\%$-$59\%$ ’yes’ answers). A manual answer analysis reveals that human annotators often decide positively despite raising several negative points. They frequently weigh the raised negative points more lightly, reserving a negative decision for extreme cases. LLMs appear to make more harsh conclusions when the accompanying textual answer contains several points of criticism. This difference in weighing weaknesses in the paper explains the comparatively low performance on decision making for the boolean nodes. Additionally, this points to the risk of a potential negativity bias of LLMs while diagnosing a paper.
Answer Language
We turn to a language analysis of the textual answers. LLMs consistently generate longer answers ($67$ tokens on average for GPT-3.5) than humans ($55$ average tokens). Upon manual inspection, we find that the automatically generated answers to extract steps tend to split up facts exhaustively, whereas humans make broader and summarizing statements. For infer steps, LLMs often restate the inputs from prior steps, unlike human annotators. Both factors might contribute to the high factuality scores by automatic metrics because LLMs appear more specific and cover many facts, whereas humans tend to select a minimal set of facts relevant to their individual assessment process. This encourages further research on automatic evaluation metrics for this setting that can account for the reasoning context of a step.
AI-Assistance Disclosure Statement
We employed LLMs to prepare this paper in a limited and specific capacity. We used LLMs to assist in revising the writing style of this document. Additionally, we employed them as assistants during the coding phase of the experiments. LLMs were not involved in the ideation process, the formulation of hypotheses, or the performance of any experiments. All scientific and intellectual contributions are solely the work of the authors.